OpenStudy (anonymous):

show that absolute value of alpha + beta is less than or equal to absolute value of alpha plus absolute value of beta for any alpha and beta. (hint f(x) as a piece wise function where alpha is 0 less than/equal to x less than or equal to 1 and beta is 1 less than or equal to beta less than or equal to 2. hint 2: use the absolute value of intergrals properties

OpenStudy (anonymous):

but how do i prove it? i need to show it.

OpenStudy (anonymous):

x less than 1 than 2 and 2 equal to

OpenStudy (anonymous):

use \[|x|\le a \implies -a \le x \le a\]

OpenStudy (anonymous):

you have change it